# Certifying Stability of Reinforcement Learning Policies using Generalized Lyapunov Functions

### v2

@@ Abstract
We study the problem of certifying the stability of closed-loop systems under control policies derived from optimal control or reinforcement learning (RL). Classical Lyapunov methods require a strict step-wise decrease in the Lyapunov function but such a certificate is difficult to construct for a learned control policy. The value function associated with an RL policy is a natural Lyapunov function candidate but it is not clear how it should be modified. To gain intuition, we first study the linear quadratic regulator (LQR) problem and make two key observations. First, a Lyapunov function can be obtained from the value function of an LQR policy by augmenting it with a residual term related to the system dynamics and stage cost. Second, the classical Lyapunov decrease requirement can be relaxed to a generalized Lyapunov condition requiring only decrease on average over multiple time steps. Using this intuition, we consider the nonlinear setting and formulate an approach to learn generalized Lyapunov functions by augmenting RL value functions with neural network residual terms. Our approach successfully certifies the stability of RL policies trained on Gymnasium and DeepMind Control benchmarks. We also extend our method to jointly train neural controllers and stability certificates using a multi-step Lyapunov loss, resulting in larger certified inner approximations of the region of attraction compared to the classical Lyapunov approach. Overall, our formulation enables stability certification for a broad class of systems with learned policies by making certificates easier to construct, thereby bridging classical control theory and modern learning-based methods.
@@ 1 Introduction
Designing and certifying stable control policies is a core challenge in both control theory and reinforcement learning (RL). Although RL methods are capable of learning policies that optimize long-term performance, their stability guarantees are often lacking. The classical Lyapunov theory (Lyapunov, 1992) offers a principled framework for certifying stability but constructing a function that satisfies the required stepwise decrease condition is often challenging in practice. While value functions from optimal control or RL encode long-term cumulative cost, they do not naturally satisfy the Lyapunov stability conditions.
@@
We propose certifying stability using a generalized notion of Lyapunov function (Fürnsinn et al., 2023, 2025) that replaces the stepwise decrease condition with a multi-step, weighted decrease criterion, allowing temporary increases over a finite horizon. As a result, constructing a generalized Lyapunov stability certificate is easier, enabling stability verification for a broader class of policies and systems.
@@
For linear systems with quadratic cost under discounted optimal control, the value function typically fails to satisfy classical Lyapunov conditions due to the influence of the discount factor. Postoyan et al. (2017) show that augmenting the value function with a quadratic residual term allows certifying stability via a linear matrix inequality (LMI), although the guarantees are conservative with respect to the discount factor. Leveraging this insight, we show that augmenting the value function with a quadratic residual leads to a valid generalized Lyapunov function, which certifies stability for a boarder range of discount factors.
@@
Building on the success of our approach in the linear setting, we extend it to nonlinear systems, where classical Lyapunov analysis is substantially more difficult. We augment value functions learned by RL algorithms (e.g., PPO (Schulman et al., 2017), SAC (Haarnoja et al., 2018)) with a neural residual term to form generalized Lyapunov functions, and jointly learn state-dependent multi-step weights to satisfy the generalized decrease condition. This joint learning scheme enables certification of nonlinear RL policies in settings where classical Lyapunov functions are challenging to construct by learning generalized certificates that tolerate non-monotonic behavior along trajectories.
@@
In addition to showing stability of fixed policies, we extend our approach to support the joint synthesis of stable neural controllers and generalized Lyapunov certificates. Prior works (Chang et al., 2019; Wu et al., 2023; Yang et al., 2024) frame this problem as learning a controller–certificate pair that maximizes the volume of a certifiable region, while iteratively training on counter-examples where Lyapunov conditions fail. We incorporate our generalized multi-step condition into this formulation and show that it enables larger formally verifiable inner approximations of the region of attraction compared to standard one-step Lyapunov training.
@@
  • We propose a new way to certify the stability of control policies obtained from optimal control or reinforcement learning by defining a generalized Lyapunov function as a combination of the policy’s value function and a neural residual term.
@@
  • We show analytically that for the linear quadratic regulator (LQR) problem, the validity of the proposed generalized Lyapunov function can be certified by a set of LMIs.
  • We show empirically that our approach allows certifying the stability of RL policies trained on Gymnasium (Towers et al., 2024) and DeepMind Control Suite (Tassa et al., 2018) benchmarks and leads to larger verified inner approximations of the region of attraction when the policy and generalized Lyapunov function are jointly optimized.
@@
Related Work. Optimal control (Bertsekas, 2012) and reinforcement learning (RL) (Sutton et al., 1998) offer a variety of principled techniques for designing control policies to optimize the performance of intelligent agents and dynamical systems. In the special case of linear systems with quadratic rewards, the optimality and stability of LQR policies (Anderson and Moore, 2007) are well understood. However, despite a variety of methods for synthesizing policies (e.g., PPO (Schulman et al., 2017), SAC (Haarnoja et al., 2018), and DDPG (Lillicrap et al., 2015) are among the most widely used) for nonlinear systems with nonlinear costs, very few techniques are available for certifying the stability of the resulting closed-loop system.
@@
Several works have explored connections between Lyapunov stability (Lyapunov, 1992) and RL, mainly in the context of safe RL. Perkins and Barto (2002) proposed a safe RL formulation by learning to switch among hand-designed Lyapunov-stable controllers. Berkenkamp et al. (2017) developed a model-based RL algorithm that uses Gaussian process dynamics and Lyapunov verification to safely improve policies while guaranteeing stability with high probability. Chow et al. (2018) formulate a Lyapunov-based method for safe RL in constrained Markov decision processes, using linear constraints to ensure safety during training. Hejase and Ozguner (2023) regularized policy training by penalizing violations of a learned Lyapunov condition to encourage stability in autonomous driving. Mittal et al. (2020) learn a Lyapunov function from suboptimal demonstrations and incorporate it as a terminal cost in a one-step model predictive control formulation to improve stability.
@@
In parallel, there has been growing interest in jointly learning policies and Lyapunov functions from data to obtain stability guarantees for nonlinear systems. Chang et al. (2019) trained Lyapunov certificates with formal verification via satisfiability modulo theories. Other works constrain neural architectures to satisfy Lyapunov conditions structurally (Gaby et al., 2022; Boffi et al., 2021), or apply mixed-integer programming (Dai et al., 2021). Long et al. (2023) studied neural distributionally robust Lyapunov functions under additive uncertainty. Wu et al. (2023) learn provably stable controllers for discrete-time nonlinear systems by jointly training a policy and Lyapunov function, with formal verification via mixed-integer linear programs. Yang et al. (2024) use gradient-guided falsification for training a controller-certificate pair, and apply formal verification via $\alpha,\beta$-CROWN (Wang et al., 2021) to certify stability post training. See also Dawson et al. (2023) for a comprehensive survey.
@@ 2 Problem Formulation
$J_{\gamma}^{*}(\mathbf{x}_{0})=\min_{\boldsymbol{\pi}}$ $\;J_{\gamma}^{\boldsymbol{\pi}}(\mathbf{x}_{0}):=\sum_{k=0}^{\infty}\gamma^{k}\,\ell(\mathbf{x}_{k},\boldsymbol{\pi}(\mathbf{x}_{k})),$ (2)
@@
Theorem 2.2 (Asymptotic Stability via a Lyapunov Function (Khalil, 1996, Theorem 3.3)).
@@ 3 Lyapunov Stability Analysis for Linear Quadratic Problems
To understand the relationship between a discounted value function $J_{\gamma}^{\boldsymbol{\pi}}$ obtained from (2) and a Lyapunov function $V$ certifying stability of the closed-loop system (3), we first study a simple setting with a linear system and quadratic stage cost. Consider the discrete-time linear system: $\mathbf{x}_{k+1}=\mathbf{A}\mathbf{x}_{k}+\mathbf{B}\mathbf{u}_{k}$, where $\mathbf{x}_{k}\in\mathbb{R}^{n}$, $\mathbf{u}_{k}\in\mathbb{R}^{m}$, and $\mathbf{A}$, $\mathbf{B}$ are known constant matrices.
Choosing the stage cost in (2) as $\ell(\mathbf{x},\mathbf{u})=\mathbf{x}^{\top}\mathbf{Q}\mathbf{x}+\mathbf{u}^{\top}\mathbf{R}\mathbf{u}$ with $\mathbf{Q}=\mathbf{C}^{\top}\mathbf{C}\succeq 0$ and $\mathbf{R}\succ 0$ leads to the discounted LQR problem (Anderson and Moore, 2007). The optimal value function $J_{\gamma}^{*}(\mathbf{x})=\mathbf{x}^{\top}\mathbf{P}_{\gamma}\mathbf{x}$ is quadratic, where $\mathbf{P}_{\gamma}$ solves the discounted Algebraic Riccati Equation (ARE):
@@
Although $J_{\gamma}^{*}$ satisfies the discounted Bellman equation, it is well understood that it does not guarantee Lyapunov decrease due to $\gamma\in(0,1)$ (De Farias and Van Roy, 2003). Moreover, verifying stability by directly analyzing the eigenvalues of $\mathbf{F}_{\gamma}^{*}$ and solving the discounted ARE (5) is nontrivial due to its nonlinear dependence on $\gamma$. This motivates the idea (Postoyan et al., 2017) of constructing valid Lyapunov functions by augmenting $J_{\gamma}^{*}$ with a residual term. To formalize this, we first state standard assumptions that ensure the existence of a stabilizing policy and a well-defined value function.
@@
Theorem 3.2 (Lyapunov Stability via LMIs for Discounted LQR (Postoyan et al., 2017)).
@@
Example 3.3 ((Postoyan et al., 2017, Example 1)).
@@ 4 Generalized Lyapunov Stability and Applications to Linear Systems
Building on the observation from Example 3.3, we introduce a generalized notion of Lyapunov stability (Fürnsinn et al., 2023, Definition 2.2) that relaxes the classical pointwise decrease condition.
@@
$\frac{1}{M}\sum_{i=1}^{M}\sigma_{i}(\mathbf{x})V(\mathbf{x}_{i})-V(\mathbf{x})<0,$ (9)
@@
See Appendix A for the proof (adapted from (Fürnsinn et al., 2023, Theorem 2)) and Appendix B for a method to construct classical Lyapunov functions from generalized ones.
@@
Note that when $\sigma_{1}=M$ and $\sigma_{i}=0$ for all $2\leq i\leq M$, Theorem 4.3 recovers the classical Lyapunov result in Theorem 3.2. Therefore, if the original one-step LMIs (7) are feasible, then feasible solutions for (10) also exist.
@@ 5 Nonlinear Systems: Stability Certification of RL Policies
Inspired by the treatment of linear systems described in Section 4, we formulate an approach to certify stability of policies for nonlinear systems obtained by RL using a generalized Lyapunov function. The key observation from Theorem 4.3 is that augmenting the optimal value function with a residual term can result in a valid generalized Lyapunov function, as in (11). Here, we use the same idea to form generalized Lyapunov functions for nonlinear systems with unknown dynamics, which is a problem considered by RL. The value function and policy obtained by an RL algorithm are typically parameterized by neural networks. Let $\boldsymbol{\pi}_{\text{RL}}$ be a pre-trained RL policy (e.g., obtained by (Schulman et al., 2017; Haarnoja et al., 2018; Hansen et al., 2022)), and let $J^{\boldsymbol{\pi}_{\text{RL}}}_{\gamma}$ denote its corresponding learned value function. We consider a generalized Lyapunov candidate as:
@@
$F(\mathbf{x}_{k}):=\frac{1}{M}\sum_{i=1}^{M}\sigma_{i}(\mathbf{x}_{k};\boldsymbol{\theta}_{2})\,V(\mathbf{x}_{k+i};\boldsymbol{\theta}_{1})-(1-\bar{\alpha})\,V(\mathbf{x}_{k};\boldsymbol{\theta}_{1}),$ (14)
@@
Training and Evaluation Setup. We evaluate our method on two standard RL control benchmarks from Gymnasium (Towers et al., 2024) and the DeepMind Control Suite (Tassa et al., 2018). RL policies $\boldsymbol{\pi}_{\text{RL}}$ and their corresponding value functions $J^{\boldsymbol{\pi}_{\text{RL}}}_{\gamma}$ are trained using implementations from Raffin et al. (2021); Hansen et al. (2022). We collect a total of $N$ rollout trajectories by simulating the closed-loop system under $\boldsymbol{\pi}_{\text{RL}}$, with each trajectory $\tau^{(i)}=\{\mathbf{x}_{k}^{(i)}\}_{k=0}^{M}$ starting from a randomly sampled initial state $\mathbf{x}_{0}^{(i)}$. Additional training and architecture details are provided in Appendix D.
Inverted Pendulum Swingup. We consider the inverted pendulum environment from Towers et al. (2024) with parameters $m=1$, $l=1$, $g=10$, and control limits $|u|\leq\frac{mgl}{5}=2$. The state space is $[-\pi,\pi)\times[-8,8]$. Due to tight torque limits, the pendulum must swing back and forth to build momentum before reaching the upright position. This makes it challenging to synthesize a policy with a Lyapunov certificate valid over the entire state space. Prior work (Yang et al., 2024; Long et al., 2024) has shown that regions where stability can be verified are typically restricted to small neighborhoods near the upright position, failing to cover the full swing-up trajectories. In contrast, modern RL policies can discover effective swing-up behaviors but lack stability guarantees. Our method bridges this gap: we take a trained SAC policy Haarnoja et al. (2018) and apply our generalized Lyapunov function training with $M=15$. Figure 5 shows the generalized Lyapunov function values along several trajectories from different initial states. As expected, the function exhibits non-monotonic behavior but shows an overall decline over the planning horizon. Figure 5 visualizes the Lyapunov function across the state space. Figure 5 plots the residual $F(\mathbf{x}_{k})$ defined in (14), verifying that the generalized decrease condition is satisfied throughout the domain.
@@
Cartpole Swingup. We consider the cartpole swing-up task from Tassa et al. (2018). The system state is represented as $[x,\cos(\theta),\sin(\theta),\dot{x},\dot{\theta}]$, where $x$ is the cart position, $\theta$ is the pole angle, and the remaining terms are their velocities. The goal is to swing the pole upright $\theta=0$ and stabilize the position at $x=0$. We use a TD-MPC policy (Hansen et al., 2022) and apply our certificate training (15) with $M=20$. Figure 6 visualize the generalized Lyapunov condition across three representative 2D slices of the state space, with the remaining two states fixed at zero. The generalized decrease condition is satisfied throughout these slices, suggesting asymptotic stability of the RL policy.
@@
In addition to qualitative visualizations, we quantitatively evaluate our learned certificates by sampling $N_{\text{test}}$ states from the full state space and checking whether (14) is satisfied. For each environment, we train certificates for multiple RL policies (the PPO policy fails to stabilize cartpole from some initial states). As shown in Table 1, the condition holds for all test states in both environments.
@@
Environment RL Methods $\boldsymbol{M}$ $\boldsymbol{N_{\text{test}}}$ Satisfying $F(\mathbf{x}_{k})\leq 0$

### v4

@@ Abstract
Establishing stability certificates for closed-loop systems under reinforcement learning (RL) policies is essential to move beyond empirical performance and offer guarantees of system behavior. Classical Lyapunov methods require a strict stepwise decrease in the Lyapunov function but such certificates are difficult to construct for learned policies. The RL value function is a natural candidate but it is not well understood how it can be adapted for this purpose. To gain intuition, we first study the linear quadratic regulator (LQR) problem and make two key observations. First, a Lyapunov function can be obtained from the value function of an LQR policy by augmenting it with a residual term related to the system dynamics and stage cost. Second, the classical Lyapunov decrease requirement can be relaxed to a generalized Lyapunov condition requiring only decrease on average over multiple time steps. Using this intuition, we consider the nonlinear setting and formulate an approach to learn generalized Lyapunov functions by augmenting RL value functions with neural network residual terms. Our approach successfully certifies the stability of RL policies trained on Gymnasium and DeepMind Control benchmarks. We also extend our method to jointly train neural controllers and stability certificates using a multi-step Lyapunov loss, resulting in larger certified inner approximations of the region of attraction compared to the classical Lyapunov approach. Overall, our formulation enables stability certification for a broad class of systems with learned policies by making certificates easier to construct, thereby bridging classical control theory and modern learning-based methods.
@@ 1 Introduction
Designing and certifying stable control policies is a core challenge in both control theory and reinforcement learning (RL). Although RL methods are capable of learning policies that optimize long-term performance, their stability guarantees are often lacking. The classical Lyapunov theory (lyapunov1992general) offers a principled framework for certifying stability but constructing a function that satisfies the required stepwise decrease condition is often challenging in practice. While value functions from optimal control or RL encode long-term cumulative cost, they do not naturally satisfy the Lyapunov stability conditions.
@@
We propose certifying stability using a generalized notion of Lyapunov function (relaxed_LF_MPC_2023; FURNSINN2025_auto) that replaces the stepwise decrease condition with a multi-step, weighted decrease criterion, allowing temporary increases over a finite horizon. As a result, constructing a generalized Lyapunov stability certificate is easier, enabling stability verification for a broader class of policies and systems.
@@
For linear systems with quadratic cost under discounted optimal control, the value function typically fails to satisfy classical Lyapunov conditions due to the influence of the discount factor. Romain_2017_TAC show that augmenting the value function with a quadratic residual term allows certifying stability via a linear matrix inequality (LMI), although the guarantees are conservative with respect to the discount factor. Leveraging this insight, we show that augmenting the value function with a quadratic residual leads to a valid generalized Lyapunov function, which certifies stability for a boarder range of discount factors.
@@
Building on the success of our approach in the linear setting, we extend it to nonlinear systems, where classical Lyapunov analysis is substantially more difficult. We augment value functions learned by RL algorithms (e.g., PPO (schulman2017proximal), SAC (haarnoja2018soft)) with a neural residual term to form generalized Lyapunov functions, and jointly learn state-dependent multi-step weights to satisfy the generalized decrease condition. This joint learning scheme enables certification of nonlinear RL policies in settings where classical Lyapunov functions are challenging to construct by learning generalized certificates that tolerate non-monotonic behavior along trajectories.
@@
In addition to showing stability of fixed policies, we extend our approach to support the joint synthesis of stable neural controllers and generalized Lyapunov certificates. Prior works (Chang2019NeuralLC; wu2023neural; yang2024lyapunovstable) frame this problem as learning a controller–certificate pair that maximizes the volume of a certifiable region, while iteratively training on counter-examples where Lyapunov conditions fail. We incorporate our generalized multi-step condition into this formulation and show that it enables larger formally verifiable inner approximations of the region of attraction compared to standard one-step Lyapunov training.
@@
  • We introduce an approach for certifying stability of reinforcement learning policies by constructing generalized Lyapunov functions, composed of the policy’s value function and a learned neural residual. Empirically, our formulation certifies policies on Gymnasium (towers2024gymnasium) and DeepMind Control Suite (tassa2018deepmind_controlsuite) benchmarks where classical single-step Lyapunov methods fail.
@@
  • We show analytically that, for the linear quadratic regulator (LQR) problem, the validity of the proposed generalized Lyapunov function can be certified by a set of LMIs.
  • We extend the formulation to jointly learn a control policy and a stability certificate using a generalized multi-step Lyapunov loss, resulting in larger certified regions of attraction compared to classical Lyapunov approaches.
  • We provide an open-source implementation of our method at https://github.com/ExistentialRobotics/Generalized_Policy_Stability.
@@
Related Work. Optimal control (bertsekas2012dynamic) and reinforcement learning (RL) (sutton1998reinforcement) offer a variety of principled techniques for designing control policies to optimize the performance of intelligent agents and dynamical systems. In the special case of linear systems with quadratic rewards, the optimality and stability of LQR policies (anderson2007optimal) are well understood. However, despite a variety of methods for synthesizing policies (e.g., PPO (schulman2017proximal), SAC (haarnoja2018soft), and DDPG (lillicrap2015continuous) are among the most widely used) for nonlinear systems with nonlinear costs, very few techniques are available for certifying the stability of the resulting closed-loop system.
@@
Several works have explored connections between Lyapunov stability (lyapunov1992general) and RL, mainly in the context of safe RL. perkins2002lyapunov proposed a safe RL formulation by learning to switch among hand-designed Lyapunov-stable controllers. berkenkamp2017safe developed a model-based RL algorithm that uses Gaussian process dynamics and Lyapunov verification to safely improve policies while guaranteeing stability with high probability. chow2018lyapunov formulate a Lyapunov-based method for safe RL in constrained Markov decision processes, using linear constraints to ensure safety during training. hejase2023lyapunov regularized policy training by penalizing violations of a learned Lyapunov condition to encourage stability in autonomous driving. mittal2020neural learn a Lyapunov function from suboptimal demonstrations and incorporate it as a terminal cost in a one-step model predictive control formulation to improve stability.
@@
In parallel, there has been growing interest in jointly learning policies and Lyapunov functions from data to obtain stability guarantees for nonlinear systems. Chang2019NeuralLC trained Lyapunov certificates with formal verification via satisfiability modulo theories. Other works constrain neural architectures to satisfy Lyapunov conditions structurally (gaby2021lyapunov_net; boffi2021learning), or apply mixed-integer programming (dai_2021_lyapunov). long2023dro_lf studied neural distributionally robust Lyapunov functions under additive uncertainty. wu2023neural learn provably stable controllers for discrete-time nonlinear systems by jointly training a policy and Lyapunov function, with formal verification via mixed-integer linear programs. yang2024lyapunovstable use gradient-guided falsification for training a controller-certificate pair, and apply formal verification via $\alpha,\beta$-CROWN (wang2021beta) to certify stability post training. See also Dawson_2022_survey for a comprehensive survey.
@@ 2 Problem Formulation
$J_{\gamma}^{*}(\mathbf{x}_{0})=\min_{\boldsymbol{\pi}}$ $\;J_{\gamma}^{\boldsymbol{\pi}}(\mathbf{x}_{0})=\sum_{k=0}^{\infty}\gamma^{k}\,\ell(\mathbf{x}_{k},\boldsymbol{\pi}(\mathbf{x}_{k})),$ (2)
@@
Theorem 2.2 (Asymptotic Stability via a Lyapunov Function (khalil1996nonlinear, Theorem 3.3)).
@@ 3 Lyapunov Stability Analysis for Linear Quadratic Problems
To understand the relationship between a discounted value function $J_{\gamma}^{\boldsymbol{\pi}}$ obtained from (2) and a Lyapunov function $V$ certifying stability of the closed-loop system (3), we first study a simple setting with a linear system and quadratic stage cost.
Consider the discrete-time linear system: $\mathbf{x}_{k+1}=\mathbf{A}\mathbf{x}_{k}+\mathbf{B}\mathbf{u}_{k}$, where $\mathbf{x}_{k}\in\mathbb{R}^{n}$, $\mathbf{u}_{k}\in\mathbb{R}^{m}$, and $\mathbf{A}$, $\mathbf{B}$ are known constant matrices.
Choosing the stage cost in (2) as $\ell(\mathbf{x},\mathbf{u})=\mathbf{x}^{\top}\mathbf{Q}\mathbf{x}+\mathbf{u}^{\top}\mathbf{R}\mathbf{u}$ with $\mathbf{Q}=\mathbf{C}^{\top}\mathbf{C}\succeq 0$ and $\mathbf{R}\succ 0$ leads to the discounted LQR problem (anderson2007optimal). The optimal value function $J_{\gamma}^{*}(\mathbf{x})=\mathbf{x}^{\top}\mathbf{P}_{\gamma}\mathbf{x}$ is quadratic, where $\mathbf{P}_{\gamma}$ solves the discounted Algebraic Riccati Equation (ARE):
@@
Although $J_{\gamma}^{*}$ satisfies the discounted Bellman equation, it is well understood that it does not guarantee Lyapunov decrease due to $\gamma\in(0,1)$ (de2003linear). Moreover, verifying stability by directly analyzing the eigenvalues of $\mathbf{F}_{\gamma}^{*}$ and solving the discounted ARE (5) is nontrivial due to its nonlinear dependence on $\gamma$. This motivates the idea (Romain_2017_TAC) of constructing valid Lyapunov functions by augmenting $J_{\gamma}^{*}$ with a residual term. To formalize this, we first state standard assumptions that ensure the existence of a stabilizing policy and a well-defined value function.
@@
Theorem 3.2 (Lyapunov Stability via LMIs for Discounted LQR (Romain_2017_TAC)).
@@
Example 3.3 ((Romain_2017_TAC, Example 1)).
@@ 4 Generalized Lyapunov Stability and Applications to Linear Systems
Building on the observation from Example 3.3, we introduce a generalized notion of Lyapunov stability (relaxed_LF_MPC_2023, Definition 2.2) that relaxes the classical pointwise decrease condition.
@@
$\frac{1}{M}\!\biggl(\sum_{i=1}^{M}\sigma_{i}(\mathbf{x})\,V(\mathbf{x}_{i})\biggr)-V(\mathbf{x})<0,$ (9)
@@
See Appendix A for the proof (adapted from (relaxed_LF_MPC_2023, Theorem 2)), and Appendix B for an explicit construction of a classical one-step Lyapunov function from a generalized Lyapunov function when the step weights $\{\sigma_{i}\}_{i=1}^{M}$ are state-independent.
Remark 4.3 (Relation to $k$-Inductive Verification).
The generalized Lyapunov condition in (9) can be viewed through the lens of $k$-step inductive verification (brain2015safety; anand2022k; wooding2024learning). In particular, by selecting weights $\sigma_{1}=\cdots=\sigma_{k-1}=0$ and $\sigma_{k}=M$, condition (9) reduces to $V(x_{t+k})\leq(1-\alpha)V(x_{t})$, which is analogous to a $k$-step Lyapunov decrease condition used in $k$-inductive reasoning. Classical $k$-induction requires a strict decrease at the final step while permitting bounded increases in intermediate steps. In contrast, Definition 4.1 enforces a decrease in a weighted average across multiple steps, with the weights $\sigma_{i}$ chosen adaptively or learned during training. This distinction provides additional flexibility, enabling certificates that may be easier to construct for nonlinear systems with learned controllers.
@@
Note that when $\sigma_{1}=M$ and $\sigma_{i}=0$ for all $2\leq i\leq M$, Theorem 4.4 recovers the classical Lyapunov result in Theorem 3.2. Therefore, if the original one-step LMIs (7) are feasible, then feasible solutions for (10) also exist. Empirically, the multi-step formulation can certify stability in additional cases where the one-step condition is infeasible, effectively enlarging the set of stabilizable systems. Formal characterization of this weaker feasibility property is left for future work.
@@ 5 Nonlinear Systems: Stability Certification of RL Policies
Inspired by the treatment of linear systems described in Section 4, we first formulate a post-hoc approach to certify stability of policies for nonlinear systems obtained by RL using a generalized Lyapunov function. The key observation from Theorem 4.4 is that augmenting the optimal value function with a residual term can result in a valid generalized Lyapunov function, as in (11). Here, we use the same idea to form generalized Lyapunov functions for nonlinear systems with unknown dynamics, which is a problem considered by RL. The value function and policy obtained by an RL algorithm are typically parameterized by neural networks. Let $\boldsymbol{\pi}_{\text{RL}}$ be a pre-trained RL policy (e.g., obtained by (schulman2017proximal; haarnoja2018soft; Hansen2022tdmpc)), and let $J^{\boldsymbol{\pi}_{\text{RL}}}_{\gamma}$ denote its corresponding learned value function. We consider a generalized Lyapunov candidate as:
@@
$F(\mathbf{x}_{k}):=\frac{1}{M}\biggl(\sum_{i=1}^{M}\sigma_{i}(\mathbf{x}_{k};\boldsymbol{\theta}_{2})\,V(\mathbf{x}_{k+i};\boldsymbol{\theta}_{1})\biggr)-(1-\bar{\alpha})\,V(\mathbf{x}_{k};\boldsymbol{\theta}_{1}),$ (14)
@@
Training and Evaluation Setup. We evaluate our method on two standard RL control benchmarks from Gymnasium (towers2024gymnasium) and the DeepMind Control Suite (tassa2018deepmind_controlsuite). RL policies $\boldsymbol{\pi}_{\text{RL}}$ and their corresponding value functions $J^{\boldsymbol{\pi}_{\text{RL}}}_{\gamma}$ are trained using implementations from stable-baselines3; Hansen2022tdmpc. We collect a total of $N$ rollout trajectories by simulating the closed-loop system under $\boldsymbol{\pi}_{\text{RL}}$, with each trajectory $\tau^{(i)}=\{\mathbf{x}_{k}^{(i)}\}_{k=0}^{M}$ starting from a randomly sampled initial state $\mathbf{x}_{0}^{(i)}$. Additional training and architecture details are provided in Appendix D.
Inverted Pendulum Swingup. We consider the inverted pendulum environment from towers2024gymnasium with parameters $m=1$, $l=1$, $g=10$, and control limits $|u|\leq\frac{mgl}{5}=2$. The state space is $\pi,\pi)\times[-8,8]$. Due to tight torque limits, the pendulum must swing back and forth to build momentum before reaching the upright position. This makes it challenging to synthesize a policy with a Lyapunov certificate valid over the entire state space. Prior work (yang2024lyapunovstable; long_2024_dr_synthesis) has shown that regions where stability can be verified are typically restricted to small neighborhoods near the upright position, failing to cover the full swing-up trajectories. In contrast, modern RL policies can discover effective swing-up behaviors but lack stability guarantees. Our method bridges this gap: we take a trained SAC policy haarnoja2018soft and apply our generalized Lyapunov function training with $M=15$. Figure 5 shows the generalized Lyapunov function values along several trajectories from different initial states. As expected, the function exhibits non-monotonic behavior but shows an overall decline over the planning horizon. Figure 5 visualizes the Lyapunov function across the state space. Figure 5 plots the residual $F(\mathbf{x}_{k})$ defined in (14), verifying that the generalized decrease condition is satisfied throughout the domain.
@@
Cartpole Swingup. We consider the cartpole swing-up task from tassa2018deepmind_controlsuite. The system state is represented as $x,\cos(\theta),\sin(\theta),\dot{x},\dot{\theta}]$, where $x$ is the cart position, $\theta$ is the pole angle, and the remaining terms are their velocities. The goal is to swing the pole upright $\theta=0$ and stabilize the position at $x=0$. We use a TD-MPC policy (Hansen2022tdmpc) and apply our certificate training (15) with $M=20$. Figure 6 visualize the generalized Lyapunov condition across three representative 2D slices of the state space, with the remaining two states fixed at zero. The generalized decrease condition is satisfied throughout these slices, suggesting asymptotic stability of the RL policy.
@@
In addition to qualitative visualizations, we quantitatively evaluate our learned certificates by sampling $N_{\text{test}}$ states from the full state space and checking whether (14) is satisfied. For each environment, we train certificates for multiple RL policies (the PPO policy fails to stabilize the cartpole from some initial states). As shown in Table 1, the condition holds for all test states in both environments.
@@
Environment RL Methods $\boldsymbol{M}$ $\boldsymbol{N_{\text{test}}}$ % Satisfying $F(\mathbf{x}_{k})\leq 0$
@@
Concentration of Learned Step-Weights.
To further analyze the learned generalized Lyapunov function, we examine the distribution of the learned step-weights across the horizon. For each trained policy, the step-weight network was evaluated over $10,000$ uniformly sampled states from the state space, and the normalized weights were averaged across samples. The weights were then grouped into five equal bins corresponding to different segments of the Lyapunov horizon. Table 2 summarizes the fraction of the total weight assigned to each bin.
Table 2: Average fraction of total step-weight assigned to each portion of the rollout horizon. Results are averaged over 10,000 test states for each policy.
Environment + Policy Steps 0–20% 20–40% 40–60% 60–80% 80–100%
Pendulum + PPO 12.3 11.4 16.7 21.8 37.8
Pendulum + SAC 10.2 13.3 17.5 22.6 36.4
Cartpole + SAC 10.8 10.5 16.2 25.6 36.1
Cartpole + TD-MPC 14.1 13.9 16.8 24.7 30.5
As shown in Table 2, the consistent concentration of weights toward the end of the horizon reveals a key strategy learned by our method. By assigning lower weights to initial steps and higher weights to later ones, the network effectively learns to tolerate initial non-monotonic transients in the Lyapunov candidate, focusing instead on ensuring a decisive, weighted decrease over the full horizon. This behavior empirically realizes the central benefit of the generalized Lyapunov condition: providing a flexible yet formal basis for certifying complex neural policies.